OpenStudy (anonymous):

Simplify the following expression. 7 - 4[5 - (3 + 4)] A 15 B -22 C -41 D -17

OpenStudy (anonymous):

solve the bracket first; 7-4[5-7] =7-4[-2] =7+8 =15